Jeremy cashed his weekly paycheck at the bank and asked for the money in $5 and $10 bills. The bank teller gave him 24 bills worth $170. How many $5 bills did the bank teller give him?


Answer by ikleyn(52784)

.
Jeremy cashed his weekly paycheck at the bank and asked for the money in $5 and $10 bills.
The bank teller gave him 24 bills worth $170. How many $5 bills did the bank teller give him?
~~~~~~~~~~~~~~~

Let x = # of $10 bills;

    y = # of  $5 bills.


Write the system of 2 equation in 2 unknown, as you read the problem

     x +  y =  24   (1)    (total bills)

   10x + 5y = 170   (2)    (total dollars)


To solve the system, divide equation (2) by 5 (both sides). The system will take the form

     x +  y =  24   (1)    

    2x +  y =  34   (3)  


Now subtract equation (1) from equation (3). You will get

     x      = 34 - 24 = 10.


Then from equation (1)

          y = 24 - 10 = 14.


ANSWER.  10 $20-bills and 14 $5-dollar bills.

Solved.

Check it on your own that my answer is correct.